### v2

# Calibrating Verbal Uncertainty as a Linear Feature to Reduce Hallucinations

Abstract
LLMs often adopt an assertive language style also when making false claims. Such “overconfident hallucinations” mislead users and erode trust. Achieving the ability to express in language the actual degree of uncertainty around a claim is therefore of great importance. We find that “verbal [^1] In this paper, we employ the term ‘verbal’ to mean ‘pertaining to words rather than meaning or substance,’ as opposed to ‘spoken rather than written’ (refer to Merriam-Webster’s definitions: https://www.merriam-webster.com/dictionary/verbal). Readers may substitute ‘verbal uncertainty’ with ‘expressed uncertainty’ throughout the text if they find it preferable. uncertainty” is governed by a single linear feature in the representation space of LLMs, and show that this has only moderate correlation with the actual “semantic uncertainty” of the model. We apply this insight and show that (1) the mismatch between semantic and verbal uncertainty is a better predictor of hallucinations than semantic uncertainty alone and (2) we can intervene on verbal uncertainty at inference time and reduce confident hallucinations on short-form answers, achieving an average relative reduction of ~30%.  [^2] The code is available at https://github.com/facebookresearch/verbal_uncertainty_feature_calibration.
[…]
zjiad@connect.ust.hk; jadeleiyu@cs.toronto.edu; {pascalefung; ncan}@meta.com
1 Introduction
Despite their remarkable capability in utilizing their internal knowledge, LLMs often suffer from hallucinations, stating or implying facts that are not supported neither by their input nor by their training data (Ji et al., 2022; Xiao and Wang, 2021; Bang et al., 2023; Xiong et al., 2023). The issue is exacerbated when models produce hallucinations using language that suggests high confidence. Such overconfidence can cause users to rely too heavily on these responses (Zhou et al., 2024a; Kim et al., 2024), possibly resulting in harm, loss of trust in the model, or both.
While enhancing a model’s ability to generate accurate knowledge is important, it is inevitable that models have knowledge gaps. In such cases, it is important for models to express uncertainty about their knowledge or altogether abstain from answering (Tomani et al., 2024; Feng et al., 2024; Zhou et al., 2023; Zhang et al., 2024b). We refer to this expression as "verbal uncertainty (VU)" (see § 2.2). When faced with questions close to their knowledge boundary, they should qualify their answers with expressions such as: “I am not sure but …”, and when the answer is squarely beyond such boundary, they should reply: “I don’t know”. However, LLMs lack reliable mechanisms to convey their intrinsic confidence in the correctness of generated content by means of the degree of doubt expressed in their outputs (Zhou et al., 2024b).
Figure 1: Illustration of our framework. We discover a linear verbal uncertainty feature (VUF) that controls model expression of uncertainty, and apply this insight to: (1) detect hallucinations arising from the miscalibration between high semantic uncertainty (SU) and low verbal uncertainty (VU); (2) mitigate hallucinations by intervening on activations along the VUF direction at inference time to make VU more aligned with model’s SU. For example, when asked "What is the 29th largest city in England?", the model initially responds with "It’s Bournemouth", exhibiting high SU and VU. By applying the VUF to intervention, we improve the VU to better align with the SU and the response becomes "Hmm, maybe Bournemouth?" – demonstrating a nuanced expression of uncertainty.
Figure 2: Evidence of verbal-semantic uncertainty miscalibration. This plot presents the Kernel Density Estimation (KDE) for samples from TriviaQA, categorized into four classes. These classes are based on the correctness of the answers generated by Llama3.1 and the consistency in abstaining. Miscalibration is indicated by high Semantic Entropy (proxy for SU) & low VU in hallucinated answers (red), and low SU & high VU in consistently abstained answers (blue).
Our work begins with the analysis of the model representation space. We show that, similarly to refusal (Arditi et al., 2024) and other behaviours (Zou et al., 2023), the degree of uncertainty expressed by a model is mediated by a single direction, which we call the “Verbal Uncertainty Feature” (VUF). Specifically, we show that the hidden states of input questions answered with low and high verbal uncertainty can be linearly separated. This allows us to use contrastive pairs (Burns et al., 2023; Panickssery et al., 2023) to identify a single difference-in-means direction that can be intervened upon to control model expression of uncertainty.
We next leverage our findings to study hallucinations through the lens of uncertainty features. As illustrated in Figure 2, we highlight the misalignment between verbal uncertainty and the uncertainty about what meaning to convey in model outputs, i.e. semantic uncertainty, as an important factor contributing to hallucinations. We propose a novel method to detect model hallucinations that incorporates both verbal and semantic uncertainty information, and show that it outperforms detection methods that rely solely on semantic uncertainty. Next, we propose a method, Mechanistic Uncertainty Calibration (MUC), that steers LLM activations using VUF to calibrate verbal uncertainty with the model semantic uncertainty.
We demonstrate that MUC effectively reduces model confident hallucinations, achieving an average relative reduction of 29.6% in short-form QA tasks. It also induces nuances expressions of uncertainty and achieves a 28.4% improvement in the alignment between verbal and semantic uncertainty.
Our main contributions are therefore threefold:
  1. 1.
    We discovered that verbal uncertainty is mediated by a single direction in representation space, i.e. a linear Verbal Uncertainty Feature (VUF) (§ 3).
  2. 2.
    We detect hallucinations arising from the misalignment between high semantic and low verbal uncertainty by integrating both types of uncertainty (§ 4.1).
  3. 3.
    We introduce Mechanistic Uncertainty Calibration (MUC), an inference-time intervention mechanism using VUFs to calibrate verbal uncertainty with semantic uncertainty, thereby mitigating hallucinations (§ 4.2).
In addition, we introduce methods to quantify verbal uncertainty and metrics that help characterize the calibration between the two kinds of uncertainty without requiring the model to output numerical confidence estimates. Overall, this work contributes to a better understanding of LLMs, shows how to reduce hallucinations and thereby make LLMs more trustworthy.
2 Background and Motivation
The miscalibration of semantic and verbal uncertainty triggers overconfident hallucinations. To bring the discussion in a quantitative framework, we introduce in this section some definitions and measures.
2.1 Semantic Uncertainty
Semantic Uncertainty (SU) refers to the intrinsic uncertainty of an agent in the semantic meaning expressed by a statement. It reflects the confidence level of a model’s prediction, focusing on its meaning and disregarding paraphrastic variations (Lin et al., 2022; Kadavath et al., 2022; Mielke et al., 2022).
We follow Kuhn et al. (2023) in operationalizing this concept in terms of Semantic Entropy that can be calculated as follows: for a given question, one first samples multiple model answers and clusters them into semantically equivalent groups, within which the meaning of every answer can be inferred from each other. Semantic entropy is then computed as the entropy of the probability distribution over semantic equivalence answer classes.
2.2 Verbal Uncertainty
Verbal Uncertainty (VU) quantifies the degree of doubt a speaker expresses about a proposition $P$, either explicitly or implicitly (e.g., "I doubt…", "Possibly…"). We formally define it as the complement of the subjective probability a listener would associate with $P$, conditioned on the utterance $U$ and contextual information $C$:
$\text{VU}(U\mid C)=1-Pr(P\mid U,C)$ (1)
In the specific case of short-form QA, this definition can be instantiated with $U$ being the answer given by an agent in response to a question $C$.
Answer $U_{1}$ is more verbally uncertain than $U_{2}$ if a listener would conclude that proposition $P$ is more probable based on answer $U_{2}$ than based on $U_{1}$:
$Pr(P|U_{2},C)>Pr(P|U_{1},C)$
where, e.g.:
  • $P$: "Bournemouth is the 29th largest city in England"
  • $C$: "What is the 29th largest city in England?"
  • $U_{1}$: "Hmm, maybe Bournemouth"
  • $U_{2}$: "Bournemouth"
We follow recent work in expression decisiveness quantification and employ "LLM-as-a-Judge" to measure VU (Yona et al., 2024; Zheng et al., 2023). Specifically, we sample multiple answers for each question and prompt an auxiliary evaluator LLM to directly assign a VU score to each answer. The VU for a question is the average of VU scores of all answers.  [^3] See Appendix 7 for the prompt used for VU estimation. This approach has been shown to produce reliable uncertainty estimates that are highly correlated with human judgments of perceived assertiveness (Yona et al., 2024; Fagen-Ulmschneider, 2023). To further validate the robustness of "LLM-as-a-Judge", we compute sentence embedding cosine similarities with predefined prototypical uncertainty expressions and find a high correlation with VU scores returned by LLMs (see Appendix 11.2 for details).
Figure 3: Visualization of verbalized certain (blue) vs. uncertain (pink) query representations exacted from selected layers of (a) Llama-3.1-8B-Instruct, (b) Mistral-7B-Instruct-v0.3, and (c) Qwen2.5-7B-Instruct on three datasets (TriviaQA, NQ-Open, PopQA). Please refer to Appendix 11 for the visualization of representations exacted from all layers.
2.3 Hallucination arise from Miscalibration between Semantic and Verbal Uncertainty
Ideally, VU should align with SU to let the model faithfully express uncertainty in the semantic meaning of its outputs. However, observations indicate that the two types of uncertainty are not always correlated, resulting in hallucinations. In this section, we quantitatively investigate and demonstrate the miscalibration between semantic and verbal uncertainty by analyzing samples from TriviaQA using Llama3.1  [^4] See Appendix 10 for examples..
Following Kossen et al. (2024); Farquhar et al. (2024), for each question, we generate a response using a low temperature (0.1) to obtain the most likely answer, and then sample multiple responses using a high temperature (1.0). We categorize the samples into two primary groups based on the VU level of the most likely answer: those that include abstentions (abstained) and those that do not (complying). We further subdivide these categories. For complying responses, we assess whether the answers are hallucinated or correct (hallucinated/correct). For abstained, we determine if the model consistently refuses (i.e., all samples), or if it complies at least once among the multiple sampled answers (“partly abstained”)  [^5] We obtain the MANOVA results which indicate significant differences in 4 groups with p-values ¡ 0.0001..
As shown in Figure 2, abstained responses have high VU, which is expected. Consistently abstained ones have low SU, but this is not a problematic mismatch, rather an artifact of using semantic entropy as a proxy for SU: these are cases where the model “knows that it does not know” and deals with them appropriately. There is however a large segment of complying answers with high SU and low VU that are hallucinations  [^6] There are outlier hallucinations with low SU. For a detailed analysis, please refer to Appendix 10.3.: this is the focus of our intervention. We show in § 4.1 that combining predictions of VU and SU helps detecting hallucination. Moreover, we show in § 4.2 that modulating VU to better reflect SU is crucial to prevent confident hallucinations and optimize the trade-off balance between hallucinations and false abstention.
2.4 Semantic Space of LLM
Recent research supports the hypothesis that language models represent features or concepts as linear directions within their activation space (Mikolov et al., 2013; Bolukbasi et al., 2016; Elhage et al., 2022; Park et al., 2023; Ferrando et al., 2024). These features include attributes such as harmlessness (Wolf et al., 2024; Arditi et al., 2024), truthfulness (Marks and Tegmark, 2023; Li et al., 2024), sentiment (Tigges et al., 2023), and language (Bricken et al., 2023). Building on these insights, we investigate the linear representation of VU, aiming to validate its representation and control its level.
3 Verbal Uncertainty Feature (VUF)
In this section we show that verbal uncertainty is mediated by a single direction.
3.1 Feature Extraction
To identify the verbal uncertainty features (VUFs) in the model’s residual stream activations, we adopt the difference-in-means technique (Belrose, 2023), which has been shown to effectively disentangle key feature information (Panickssery et al., 2023; Marks and Tegmark, 2023; Arditi et al., 2024; Yu et al., 2024b).
We first collect a set of pairs of questions and generated answers for which the model generates answers with high VU, denoted as question $x\in\mathcal{D}_{uncertain}$, and another set for questions on which the model generates certain answers with low VU, denoted as $x\in\mathcal{D}_{certain}$. We construct $\mathcal{D}_{uncertain}$ and $\mathcal{D}_{certain}$ by choosing top $N_{uncertain}$ and bottom $N_{certain}$ pairs sorted by their VU score, which is calculated using the LLM-as-a-Judge method described in § 2.2 with the Llama3.1-70B-Instruct model (Dubey et al., 2024). To generate answers, we use the “uncertainty” eliciting prompt provided in Appendix 7.1. We then calculate the difference between the model’s mean last-token residual stream activations $\mathbf{h}^{(l)}(x)$ for each layer $l$ when processing these two sets of questions. This difference is then L2 normalized, as illustrated in the following equation:
$\begin{split}\hat{\mathbf{r}}_{\text{VU}}^{(l)}=&\,\frac{1}{|\mathcal{D}_{\text{uncertain}}|}\sum_{x\in\mathcal{D}_{\text{uncertain}}}\mathbf{h}^{(l)}(x)\\ &-\frac{1}{|\mathcal{D}_{\text{certain}}|}\sum_{x\in\mathcal{D}_{\text{certain}}}\mathbf{h}^{(l)}(x)\end{split}$ (2)
$\mathbf{r}_{\text{VU}}^{(l)}=\frac{\hat{\mathbf{r}}_{\text{VU}}^{(l)}}{\|\hat{\mathbf{r}}_{\text{VU}}^{(l)}\|}$ (3)
Figure 4: Compare VUFs exacted from different datasets from Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, and Qwen2.5-7B-Instruct
Figure 5: Mean model-generated answer verbal uncertainty on three QA datasets with varying degrees of inference-time VUF intervention (modulated by the intervention intensity $\alpha$).
3.2 Discovery of Linear Verbal Uncertainty Features
To empirically demonstrate the VUFs explained above, we adopt three closed-book short-form QA datasets: TriviaQA (Joshi et al., 2017), NQ-Open  (Kwiatkowski et al., 2019), and PopQA (Mallen et al., 2022); and consider the following models: Llama3.1-8B-Instruct (Dubey et al., 2024), Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), and Qwen2.5-7B-Instruct (Yang et al., 2024)   [^7] For simplicity, we will refer to them as Llama3.1, Mistral, and Qwen2.5, respectively..
Visualization
We extract the activations of the last token for each question at each layer from $\mathcal{D}_{uncertain}$ and $\mathcal{D}_{certain}$ and project them into a 2D space using PCA. As illustrated in Figure 3, the activations of samples labeled as "certain" and "uncertain" are clearly linearly separated into two distinct clusters, starting from the middle layers. This strongly indicates that $\mathbf{r}_{\text{VU}}^{(l)}$ represents a meaningful linear direction that reflects the VU level of questions in hidden states. We refer to $\mathbf{r}_{\text{VU}}^{(l)}$ as VUFs.
Effective Layer Selection
To identify the effective layers of VUFs, we analyze VUFs obtained from each layer of three different models. We measure the cosine similarity of distinct VUFs extracted from TriviaQA, NQ-Open, and PopQA datasets respectively. The results presented in Figure 4 show a high cosine similarity between VUFs from different datasets, particularly in the middle and subsequent layers. This pattern is aligned with visualization and consistent across all models and datasets we examined. Observations from both visualization and similarity across datasets indicate that reliable VUFs are best extracted from the middle to the last layers.
Causal Validation
We validate the causal connection between VUFs and the model’s VU by analyzing the generation behavior as we modulate the strength of the corresponding feature through simple linear interventions. Inspired by Li et al. (2024), we intervene on model activations of all tokens by steering them along a set of VUF directions.
For each layer $l$, we extract VUs $\mathbf{r}_{\text{VU}}^{(l)}\in\mathbb{R}^{d_{model}}$. Specifically, the VU feature vector $\mathbf{r}_{\text{VU}}^{(l)}$ serves as a directional guide for steering activations, as described in the equation below:
$h^{(l)}(x)\leftarrow h^{(l)}(x)+\alpha*\mathbf{r}_{\text{VU}}^{(l)}$ (4)
where $\alpha$ is the intensity of intervention, and $\mathbf{r}_{\text{VU}}^{(l)}$ is the verbal uncertainty feature at layer $l$. The results presented in Figure 5 show that adding VUFs to model activations ($\alpha>0$) increases the VU of the model outputs. Conversely, removing VUFs from activations ($\alpha<0$) decreases this uncertainty. As the intensity of VUFs ($|\alpha|$) gets stronger, the VU scores exhibit greater changes. This trend remains consistent across all models and datasets we studied. This show the potential of VU calibration in model generation. We will further explore how to leverage this phenomenon in §4.2.
Interestingly, although Qwen2.5 exhibits a similar trend, it is significantly less sensitive compared to Llama3.1 and Mistral. This is due to the normalization of the VUF. Qwen embeddings have larger norms, resulting in longer distances between clusters.
To address potential circularity concerns due to the double use of LLM in VUF extraction and VU evaluation, we validate our findings using an alternative VUF extraction method, detailed in Appendix 11.5.
VUFs are Consistent Across Different Datasets
To investigate the generalization of VUF across datasets, we use VUFs extracted from TriviaQA to control the VU level of other datasets: NQ-Open and PopQA. As shown in Figure 6, adding or removing TriviaQA VUFs increases or decreases the VU of model outputs for these datasets. These two findings indicate that VUFs are consistent across different datasets, suggesting that a universal VUF can be derived and utilized in our experiments further in the paper. Similar results using other VU scores are provided in Appendix 11.4.
Therefore, once we identify the appropriate layers for each model, these selections remain consistent across different datasets, eliminating the need to repeat the selection process for other datasets.
Figure 6: Causal Validation on NQ-Open and PopQA with the VUF exacted from the OOD dataset TriviaQA.
4 Verbal Uncertainty and Hallucination
Confident hallucinations arise from a miscalibration between VU and SU, where the model fails to express its high uncertainty in its generated output. Taking advantage of this miscalibration, we can detect hallucinations (§ 4.1). Furthermore, we mitigate confident hallucinations by calibrating these two uncertainties using VUFs discovered in § 3 (§ 4.2).
4.1 Hallucination Detection with Semantic and Verbal Uncertainty
{NiceTabular}
rrcc|cc|cc Llama Mistral Qwen
Dataset Feature AUROC ACC AUROC ACC AUROC ACC
Semantic 79.21 81.1 71.51 74.6 72.18 72.0
Verbal 72.1 80.1 68.20 72.4 72.8 72.7
TriviaQA Combined 79.71 80.8 72.99 74.6 74.71 72.7
Semantic 65.29 70.7 64.47 62.1 56.74 53.6
Verbal 54.04 71.2 65.26 62.5 61.85 60.7
NQ-Open Combined 66.02 70.3 68.96 64.8 62.36 58.7
Semantic 71.16 81.2 66.03 71.1 53.44 71.7
Verbal 62.30 81.0 71.13 71.7 75.66 76.2
PopQA Combined 75.66 81.1 73.82 75.7 75.43 76.0
Table 1: Detection Results based on Uncertainty for Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, and Qwen2.5-7B-Instruct
Dataset SEP EigenScore Our method
Calculated Probe-Predicted
AUROC ACC AUROC ACC AUROC ACC AUROC ACC
TriviaQA 66.85 66.0 64.83 53.5 79.71 80.8 73.53 80.1
NQ-Open 54.07 53.9 56.29 49.3 66.02 70.3 57.15 71.3
PopQA 70.17 65.6 59.33 44.9 75.66 81.1 74.76 81.0
Table 2: Detection Results on Llama-3.1-8B-Instruct. ‘Calculated’ means that the SE feature is computed after sampling multiple answers, ’Probe-Predicted’ means that SE is as predicted by a probe that takes as input the embeddings of the last token of the question, and therefore does not require sampling.
We propose to detect hallucinations by leveraging both verbal and semantic uncertainties. Our approach utilizes a simple logistic regression model to predict the presence of hallucinations. We demonstrate that combining VU with SU significantly enhances the detection performance.
Measuring Semantic Entropy, our proxy for SU, requires generating multiple samples and running auxiliary models (Farquhar et al., 2024). We therefore also consider training Uncertainty Probes for uncertainty quantification following (Kossen et al., 2024) to ensure cost-efficiency. These probes are linear models trained on the hidden states of LLMs to predict numerical uncertainty values. The hidden states are extracted from the last token of the question and sourced from multiple layers within the LLM  [^8] Please refer to Appendix 12.1 for implement details.. During testing, the input to the logistic regression model consists of predicted verbal and semantic uncertainties obtained from two regressor probes. [^9] Probes trained as binary classifiers over thresholded continuous values perform similarly, see Appendix 12.2.
Dataset Hallucination Rate↓ Correct. Rate↑ Refusal Rate VU/SU Disagree. Rate ↓ Correlation↑ VU for Incorrect ↑ VU for Correct
Conf./Overall Conf. Overall
before after after before after before after before after before after before after before after
Llama3.1-8B
TriviaQA 23.3 19.0 21.2 71.3 70.6 5.4 8.2 21.50 21.40 0.59 0.63 0.50 0.55 0.16 0.16
NQ-Open 40.2 26.2 32.7 50.7 47.7 9.1 19.6 35.10 18.90 0.38 0.69 0.37 0.54 0.17 0.24
PopQA 33.7 21.6 23.2 23.5 21.0 42.8 55.8 50.70 44.70 0.05 0.34 0.61 0.73 0.17 0.20
Average 32.4 22.3 25.7 48.5 46.4 19.1 27.9 35.80 28.30 0.34 0.55 0.49 0.61 0.17 0.20
Mistral-7B
TriviaQA 30.2 19.7 26.8 67.9 67.0 1.9 6.2 27.50 16.80 0.46 0.66 0.19 0.39 0.04 0.05
NQ-Open 52.2 40.8 46.9 41.7 39.4 6.1 13.7 46.80 19.80 0.24 0.58 0.23 0.40 0.07 0.10
PopQA 58.2 26.7 32.5 26.4 23.9 15.4 43.6 50.80 28.50 0.15 0.53 0.30 0.64 0.07 0.15
Average 46.9 29.1 35.4 45.3 43.4 7.8 21.2 41.70 21.70 0.28 0.60 0.20 0.50 0.06 0.10
Qwen2.5-7B
TriviaQA 37.9 23.4 34.4 58.6 58.1 3.5 7.5 27.10 22.20 0.57 0.59 0.43 0.51 0.14 0.14
NQ-Open 61.6 46.8 56.5 30.4 30.1 8.0 13.4 44.50 32.50 0.31 0.38 0.39 0.46 0.18 0.19
PopQA 44.8 33.6 38.3 18.1 16.4 37.1 45.3 46.80 43.00 0.08 0.22 0.69 0.75 0.21 0.21
Average 48.1 34.6 43.1 35.7 34.9 16.2 22.1 39.50 32.60 0.32 0.39 0.51 0.57 0.17 0.18
Llama3.1-70B
TriviaQA 12.1 10.1 11.8 87.0 86.8 0.9 1.4 7.5 7.1 0.71 0.80 0.29 0.35 0.06 0.07
NQ-Open 35.7 32.3 34.0 60.8 59.5 3.5 6.5 21.1 15.1 0.49 0.73 0.27 0.36 0.08 0.09
PopQA 41.4 28.0 35.2 44.6 42.4 14.0 22.4 22.2 14.8 0.59 0.75 0.48 0.62 0.17 0.18
Average 29.7 23.5 27.0 64.1 62.9 6.1 10.1 16.9 12.3 0.60 0.76 0.35 0.44 0.10 0.11
Table 3: Mitigation Results for Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, Qwen2.5-7B-Instruct, and Llama-3.1-70B-Instruct. ‘Before’ represents the original generation and ‘after’ represents the generation after Mechanistic Uncertainty Calibration. The original generation is always confident, so there is no difference between ‘Confident’ and ‘Overall’.
To evaluate the performance of hallucination detection, we follow Kossen et al. (2024); Orgad et al. (2024) and adopt the area under the receiver operating characteristic curve (AUROC) as the main metric. Additionally, we use accuracy (ACC) as a reference metric.
Baselines
We adapt SEP (Kossen et al., 2024), a hallucination detection method that trains a probe to predict binarized SU based on hidden states. We employ the sentence-form and TBG (token before generating) settings of SEP. In contrast to the original setup, we classify abstained samples as non-hallucinated.
Result
As shown in Tab. 4.1, incorporating VU alongside SU improves detection performance for all models. The accuracy when using probe-predicted uncertainties is similar to that obtained when using calculated SU (Tab. 4.1). This is important because it means that it is possible to predict a high risk of hallucination already after the prefill stage of decoding, before starting autoregressive generation.
4.2 Hallucination Mitigation via Inference-time Mechanistic Uncertainty Calibration
In § 3, we observed the existence of universal VUFs extracted from the middle layers to the last. These VUFs enable us to modulate the degree of verbal uncertainty in model responses. Building on the insights, we propose Mechanistic Uncertainty Calibration (MUC). This method leverages VUFs to calibrate verbal uncertainty with semantic uncertainty.
For each layer $l$, we extract VUF vectors from the last token of question, $\mathbf{r}_{\text{VU}}^{(l)}\in\mathbb{R}^{d_{model}}$. We then modulate the influence of these features through straightforward linear interventions on all tokens in detectd hallucinated responses. Specifically, $\mathbf{r}_{\text{VU}}^{(l)}$ serves as a directional guide for steering activations, as described in the equation:
$h^{(l)}(x)\leftarrow h^{(l)}(x)+\alpha_{\text{su}}(x)*\mathbf{r}_{\text{VU}}^{(l)}$ (5)
where the magnitude of intervention
$\alpha_{\text{su}}(x)=clip(su(x)_{norm}-vu(x),0,max_{\alpha})$ (6)
is the gap between min-max normalized SU and VU of input questions x, and $\mathbf{r}_{\text{VU}}^{(l)}$ is the VUF at layer $l$. Please refer to Appendix 13.1 for details.
Unlike previous work on abstention (Feng et al., 2024; Zhang et al., 2024b), i.e. which advocate declining to answer in the face of uncertainty, we aim to incorporate nuanced uncertainty in the output text form. Our method does not involve fine-tuning of the model, additional system prompt design, or sampling methods required by previous mitigation works on nuanced uncertainty. We show the existence of universal VUFs that can be pre-comuputed, which saves on computation overhead. Our method leverages the model’s underutilized inherent ability to express nuanced uncertainty, allowing it to better manage and communicate its confidence levels.
Evaluation Metrics
To evaluate the level of hallucination and the calibration of verbal and semantic uncertainties in generated responses, we use the following metrics:
  • Overall Hallucination Rate: The proportion of samples where the model doesn’t refuse and the answer is not entailed by the golden answer.
  • Confident Hallucination Rate: The proportion of responses that are not entailed by the golden answer and have a low VU below a predefined threshold. The threshold is identified by minimizing the sum of squared distances from VU to the threshold (Kossen et al., 2024) [^10] In the original output, confident hallucinations (incorrect and certain) and hallucinations (incorrect and non-refusal) are the same, as verbally uncertain answers without refusal are rare.
  • Correctness Rate: The proportion of samples entailed by the golden answer.
  • Refusal Rate: The proportion of samples where the model refuses to answer the question.
  • VU/SU Disagreement Rate: The proportion of samples where SU and VU disagree, meaning one is above the threshold while the other is below. A lower disagreement rate suggests that the two uncertainties are well-calibrated.
  • Correlation Coefficient: The correlation coefficient between SU and VU measures the strength and direction of the linear relationship between these two uncertainties. A higher correlation coefficient indicates better alignment.
  • VU for Incorrect answer: The average of VU for incorrect responses. VU should be relatively high, indicating that the model is less confident in its incorrect outputs.
  • VU for Correct answer: The average of VU for correct responses. This serves as a reference metric to ensure that the VU for correct answers is relatively stable.
Setting
Conf. Hallu.
Rate ↓
Disagree.
Rate ↓
Corr.↑
VU for
Incorrect↑
VU for
correct
TriviaQA
w/ calculated Us 19.0 21.4 0.63 0.55 0.16
w/ predicted Us 22.3 13.5 0.86 0.49 0.20
NQ-Open
w/ calculated Us 26.2 18.9 0.69 0.54 0.24
w/ predicted Us 28.5 25.1 0.65 0.48 0.26
PopQA
w/ calculated Us 21.6 44.7 0.34 0.73 0.20
w/ predicted Us 29.7 42.3 0.41 0.59 0.39
Table 4: Ablation Study Results for Llama-3.1-8B-Instruct, showing the impact of replacing calculated uncertainty values with values predicted by probes on the hidden state of the last token of the question.
Setting
Conf. Hallu.
Rate ↓
Disagree.
Rate ↓
Corr.↑
VU for
Incorrect↑
VU for
correct
TriviaQA
before 23.3 21.5 0.59 0.5 0.16
w/ Rand 20.1 22.4 0.59 0.5 0.17
w/ VUF 19.0 21.4 0.63 0.55 0.16
NQ-Open
before 40.2 35.1 0.38 0.37 0.17
w/ Rand 35.2 26.7 0.45 0.38 0.17
w/ TriviaQA VUF 26.2 19.6 0.70 0.54 0.24
w/ VUF 26.2 18.9 0.69 0.54 0.24
PopQA
before 33.7 50.7 0.05 0.61 0.17
w/ Rand 28.8 47.6 0.16 0.63 0.18
w/ TriviaQA VUF 22.4 40.2 0.37 0.70 0.23
w/ VUF 21.6 44.7 0.34 0.73 0.20
Table 5: Ablation Study Results for Llama-3.1-8B-Instruct when the VUF from TriviaQA, NQ-Open, and PopQA datasets is replaced with: (1) VUF extracted from TriviaQA only, applied to intervene on NQ-Open and PopQA samples. (2) random vectors, applied to intervene on three datasets.
Result
We compare results before and after the application of MUC with calculated uncertainties on TriviaQA, NQ-Open, and PopQA in Tab. 3. Overall, the application of MUC results in a substantial reduction of confident hallucinations at the cost of a small decrease in Correctness Rate.  [^11] The correctness rate decreases when the correct answer exhibits high SU and the MUC intervention results in an abstention. This reduction stems from noise in measuring SU. For the same question, the sampled answers may contain diverse additional information, leading to a high calculated semantic entropy. Please see Appendix 13.3 for an example. Different trade offs between Hallucination Rate and Correct Rate can be obtained by varying the strength of the inference-time interventions. The optimal trade-off is highly application-dependent. The observed decrease in the VU/SU Disagreement Rate, along with an increase in the Correlation Coefficient, demonstrates improved calibration and alignment between verbal and semantic uncertainties across most models and datasets. While the VU for incorrect answers increased significantly, indicating reduced confidence in incorrect outputs, the VU for correct answers remained relatively unchanged after uncertainty calibration. [^12] The increase in VU also for correct answers is not necessarily wrong, as it stems from the cases where the answer is correct, but the model is semantically uncertain about it. The fact that the trend is consistent across different models and sizes underscores the generality and effectiveness of our approach.
As demonstrated in Tab. 5, mitigation results using probe-predicted uncertainties are somewhat worse but comparable to those based on calculated uncertainties. This suggests that probes can effectively predict uncertainties and mitigate hallucinations during inference.
To demonstrate the generality of our method across datasets, we applied VUFs derived from TriviaQA with calculated uncertainties to mitigate hallucinations in NQ-Open and PopQA. As illustrated in Tab. 5, the hallucination rate decreases, supporting the finding in § 3.2 that VUFs are consistent across different datasets. VUFs from one dataset can effectively control the VU level in other datasets.
To further prove the importance of using VUF in the MUC method, we perturb activations with random vectors with the same $\alpha$ and value range as the VUF, when the hallucination detector trigger. As shown in Tab. 5, applying random perturbations upon hallucination detection does lead to an improvement over the baseline. However, the results show that intervening on the VUF direction is significantly more effective than random perturbations.
This approach not only mitigates hallucinations but also enhances the overall reliability of the model’s outputs.
5 Related Work
We discuss relevant work on linear feature discovery and model steering in § 2.4. Here we present related work on other aspects of this work.
5.1 Uncertainty in LLMs
Uncertainty estimation has long been a cornerstone of reliable machine learning systems (Gawlikowski et al., 2023). Recent success of LLMs in real-world applications have expanded the scope of uncertainty estimation research, prompting efforts to address unique challenges these models face in open-ended generation (Huang et al., 2024; Duan et al., 2024). One line of research focuses on token-level uncertainty estimation over probabilities of LLM generations, such as predictive confidence or entropy. However, these approaches work at the level of tokens and do not capture the uncertainty over semantic meaning. Resampling-based methods address this limitation leveraging self-consistency across multiple generated responses (Duan et al., 2024; Zhang et al., 2024a; Farquhar et al., 2024; Wang et al., ; Malinin and Gales, 2020; Chen et al., 2024; Gao et al., 2024). However, these methods do not fully account for the verbal uncertainty expressed by models. Mielke et al. (2022) refers to a model’s verbalized expression of confidence as "linguistic confidence". This study involves manually annotating responses with labels of refusal, low or high confidence and training a specialized classifier on this dataset. Similarly, Tomani et al. (2024) introduces the concept of "in-dialogue uncertainty" by counting hedge words from a predefined list in model outputs.
5.2 Hallucination Detection and Mitigation
Detection
Studies have demonstrated that model uncertainty can serve as an indicator for identifying hallucinations (Farquhar et al., 2024; Chen et al., 2024; Zhang et al., 2023; Xiao and Wang, 2021). Other works have explored using the internal states of LLMs for detection (Azaria and Mitchell, 2023; Ji et al., ; Snyder et al., 2023; Kadavath et al., 2022). Additionally, some studies have focused on building annotated datasets and fine-tuning hallucination detectors on them (Ji et al., 2024; Gu et al., 2024; Mishra et al., 2024; Li et al., 2023a; Muhlgay et al., 2024; Varshney et al., 2023; Yang et al., 2023). To the best of our knowledge, ours is the first work to show the effectiveness of combining VU and SU for hallucination detection.
Mitigation
One approach to mitigating hallucinations is generating more faithful and factual answers include model editing (Daheim et al., 2023; Ji et al., 2023a), decoding rectification (Rebuffel et al., 2022; Chuang et al., 2023; Shi et al., 2023; Li et al., 2023b), mechanistic fine-tuning (Yu et al., 2024a; Wu et al., 2024), re-ranking (Gu et al., 2024) and variants of the Chain-of-Thought approach involving verification or reflection (Dhuliawala et al., 2023; Lei et al., 2023; Ji et al., 2023b; Wang et al., 2023). Alternative mitigation methods for improving the trustworthiness of models involve the use of abstention and controlled stopping mechanisms  (Cheng et al., 2024; Duan et al., 2024; Tomani et al., 2024; Feng et al., 2024; Zhang et al., 2024b). These works aim to completely refrain from answering the question when the model is uncertain, thereby reducing the likelihood of hallucinations.
Unlike abstention, which involves refusing to answer in the face of uncertainty, we aim to incorporate the uncertainty in the output text form. With similar motivation to this work, Band et al. (2024) trains models to verbally convey the probability that their claims are correct; Stengel-Eskin et al. (2024) fine-tunes the model based on user feedback regarding the perceived correctness of answers. Our work does not involve fine-tuning of the LLM, additional system prompt design, or sampling methods required by previous mitigation works.
6 Conclusion and Future Work
We address the critical issue of hallucinations with overconfidence in LLMs. We demonstrate that an underlying issue contributing to hallucinations is the misalignment between models’ intrinsic semantic uncertainty and its expressed verbal uncertainty (VU). We discover the existence of a Verbal Uncertainty Feature (VUF), a single direction in the representation space that governs the VU. We leverage these insights in two applications: (1) a hallucination detection method that integrates semantic and verbal uncertainties, outperforming methods relying solely on semantic uncertainty; (2) a mitigation method, Mechanistic Uncertainty Calibration (MUC), based on aligning verbal uncertainty with the model’s semantic uncertainty by steering activations along the VUF direction during inference. Our findings suggest that LLMs can benefit from a more nuanced expression of uncertainty. Empirical results demonstrate a significant reduction in hallucinations and improved alignment, thereby enhancing the trustworthiness and reliability of LLM outputs.
For future work, one potential direction is to improve the generalizability of VUF across different LLM architectures and extend its application to long-form QA tasks. Additionally, investigating why activation steering sensitivity varies among LLM architectures could provide valuable insights. Finally, exploring how models represent uncertainty arising from various causes, such as underspecified questions, controversial topics, ethical dilemmas, and model ignorance, etc., would be an interesting next step in interpretability.
Limitations
While our discovery of the Verbal and Semantic Uncertainty Framework (VUF) and the proposed method show promise in reducing hallucinations by calibrating uncertainty, there are several limitations to consider. Firstly, our investigation is based on short-form (sentence-length) QA datasets, which may not fully capture the complexity of real-world scenarios. Additionally, although we have demonstrated improvements in uncertainty calibration, the method’s reliance on predefined probes and scores may limit its adaptability to unforeseen contexts or novel queries. Lastly, our approach primarily focuses on enhancing the model’s internal mechanisms for expressing uncertainty, which does not necessarily lead to correcting hallucinated answers. Future work should address these aspects to develop a more comprehensive solution for mitigating hallucinations in LLMs.